OpenStudy (anonymous):

Hey guys i need some help with this one, The function is f(x) = 9x + 7 Then f (f(2)) = ???????

OpenStudy (anonymous):

It looks like you need to take 2 and plug it into your original equation, f(2)=9(2) +7= y then take your answer y and plug it into the original equation f(y)=9(y)+7= ? You should be able to figure the answer from here.

OpenStudy (anonymous):

ok awesome! Thanks!